### v1

# Model-based Manipulation of Deformable Objects with Non-negligible Dynamics as Shape Regulation

Abstract
Model-based manipulation of deformable objects has traditionally dealt with objects in the quasi-static regimes, either because they are extremely lightweight/small or constrained to move very slowly. On the contrary, soft robotic research has made considerable strides toward general modeling and control - despite soft robots and deformable linear objects being very similar from a mechanical standpoint. In this work, we leverage these recent results to develop a fully dynamic framework of slender deformable objects grasped at one of their ends by a robotic manipulator. We introduce a dynamic model of this system using functional strain parameterizations and describe the manipulation challenge as a regulation control problem. This enables us to define a fully model-based control architecture, for which we can prove analytically closed-loop stability and provide sufficient conditions for steady state convergence to the desired manipulation state. The nature of this work is intended to be markedly experimental. We propose an extensive experimental validation of the proposed ideas. For that, we use a 7-DoF robot tasked with the goal of positioning the distal end of six different electric cables, moving on a plane, in a given position and orientation in space.
Note: a video showcasing the performance of the proposed modeling and control strategies is available here.
I Introduction
Deformable objects are ubiquitous in our daily lives, but robots still find it extremely complex to manipulate them [1, 2, 3, 4]. A central reason for this difficulty is the need for infinite degrees of freedom (DoFs) to fully describe a deformable object’s state, making the direct application of established strategies unfeasible [5, 6, 7, 8].
This is the case also for the simplest object geometries that we will be the focus of the present work: slender or deformable linear [^1] It is worth stressing that linear does not refer to the dynamics or kinematics of these systems being linear, but to the fact that these objects look like a curved line. In fact, their dynamics is highly nonlinear, as we will discuss later in the paper. To avoid confusion, we sometimes employ the less common denomination of slender object. objects (DLOs). These are deformable bodies whose spatial configuration can be largely specified in one dominant spatial dimension. A substantial body of literature focuses on very lightweight, small, or slowly-moving DLOs (i.e., quasi-static regimes). These assumptions allow circumventing the challenge of formulating their dynamics by focusing on purely kinematic descriptions [9, 10, 11, 12, 13], usually relying on the deformation Jacobian paradigm and shape servoing. Machine learning-based variations have also been considered [14, 12]. Learning strategies are nowadays a common choice when it comes to relaxing the quasi-static hypothesis to move into the dynamic manipulation realm [15, 16, 17, 18, 19, 20, 21, 22]. However, learning-based control comes with its own well-known limitations. To name a few: lack of guarantees, need for large amounts of data, and potential unsafety of the learning process.
Figure 1: Pictorial representation of the task that we investigate in this work: the manipulation of slender deformable objects via a generic manipulator holding them to one of their ends by its end effector. We superimpose the reference frames important to describe the dynamics of the slender object.
This work aims to introduce and experimentally validate a fully model-based perspective on manipulating slender deformable objects that does not require applying any quasi-static assumption on the object’s behavior. To achieve this goal, we put ourselves in the context of classical control of deformable mechanical systems [23, 24, 25, 26, 27]. More precisely, we borrow from recent advancements in model-based control of soft robots [28], which are systems entirely made of continuously deformable soft materials [29, 30]. The past few years have seen an exponential increase in control-oriented models of soft robots [31, 32]. These include shape [^2] Their kinematic version have been independently investigated in deformable objects literature [33, 34, 35]. [36, 37] and - more recently - strain parametrizations [38, 39, 40, 41]. In turn, the availability of finite dimensional interpretable models has driven the flourishing of model-based control techniques [42, 43, 44, 45, 46, 47, 48, 49, 50, 51] that combine improved performance with theoretical guarantees.
To conclude, we propose a dynamic modeling framework for DLOs that have been pre-grasped at one of their ends by a generic robotic manipulator (see Fig. 1). The framework uses strain-based functional parameterization to model DLOs with a non-negligible dynamic response. We then cast the manipulation challenge as a regulation control problem and introduce a control architecture (see Fig. 2), for which we analytically prove closed-loop stability.
Finally, our principal goal is to showcase the possibilities that pursuing a model-based control route can yield. So, the key feature of this work is to extensively validate the proposed model in static and dynamic regimes and the control strategy when applied to object endpoint position and orientation regulation. Our setup (see Fig. 4) comprises a 7-DoF robot manipulator that grasps high-voltage electric cables normally used for charging electric vehicles.
In a broad sense, this paper can be considered the journal extension of the conference paper [52]. We preliminarily explored these ideas there, even if in a substantially simplified context and with different methods and goals.
II Strain-Based Dynamic Modeling of DLOs
Our goal is to introduce a fully dynamic model of a robot-object pair as described by Fig. 1. We start by introducing a dynamic model of the object - as it is the part that requires most of the attention - and only at the end we combine it with the model of the robot.
Compared to the model of a soft manipulator, the object part of the dynamic model introduced in this section presents the extra complexity of not having one of its ends rigidly fixed to the ground. We can look at this as if the reference frame $\{S\}$ in Fig. 1 is a floating base of the object. We also define a fixed frame representing the manipulator base, $\{S_{\mathrm{B}}\}$, and the three floating base configuration variables $x\in\mathbb{R}$, $y\in\mathbb{R}$ and $\phi\in[-\pi,\pi)$. We consider the nominal state of the object to be hanging vertically downwards from its base, so it is convenient to define an intermediate frame $\{S_{B^{\prime}}\}$ coincident with $\{S_{\mathrm{B}}\}$ but rotated $\pi$ radians in the $xy$ plane. When $x$, $y$ and $\phi$ are all 0, the object base frame $\{S\}$ is coincident with $\{S_{B^{\prime}}\}$.
II-A Additional assumptions
Here, we consider a few assumptions for the sake of simplicity of derivations but without loss of generality. We considered the object inextensible and not experiencing shear strain. We focus on the polynomial curvature functional parameterization [53], which is representative of the general strain case discussed in [39]. Moreover, we propose derivations in the case of planar motions of the object, as this is the case we consider in our experimental validation, and - again - including full motions in 3D only makes derivations more complex without any substantial shift in the core ideas introduced here [54, 55]. Furthermore, the control algorithms that we propose later in this manuscript are agnostic to these simplifying assumptions and can be applied to the general case without modifications.
II-B Object’s Backbone Kinematics
We describe the DLO shape through its backbone (i.e., its central axis) as the spatial curve $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}}):[0,1]\rightarrow\mathrm{SO}(2)\simeq\mathbb{R}^{3}$, where $s\in[0,1]$ is the local abscissa along the object. For example, $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(0)$ represents the configuration of the grasped location, $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(1)$ of the one of the other end, and $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(0.5)$ the one of the mid-point.
Using standard results in differential geometry of curves [56], we express the orientation of each point of the backbone as
$\alpha_{\mathrm{B}}(s)=(\pi+\phi)+\int_{0}^{s}\kappa(v)\mathrm{d}v,$ (1)
where $\kappa:[0,1]\rightarrow\mathbb{R}$ is the curvature strain (well-defined thanks to the Lie structure of $\mathrm{SO}(2)$), and $\phi\in S^{1}\simeq[-\pi,\pi)$ is the orientation of the grasping location.
Again, relying on the differential geometry of curves [56], the Cartesian coordinates of the central axis are evaluated by integrating components in the $x$- and $y$-directions of $\{S_{v}\}$ as $\alpha_{\mathrm{B}}$ varies up to $s$
$\begin{bmatrix}x_{\mathrm{B}}(s)\\ y_{\mathrm{B}}(s)\end{bmatrix}=\begin{bmatrix}x\\ y\end{bmatrix}+L\int_{0}^{s}\begin{bmatrix}\sin(\alpha_{\mathrm{B}}(v))\\ -\cos(\alpha_{\mathrm{B}}(v))\end{bmatrix}\mathrm{d}v$ (2)
where $L$ is the total length of the object, and $x,y\in\mathbb{R}$ are two extra parameters. It is immediate to verify that $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(0)=(x,y,\phi)$ thus giving an immediate interpretation to the three parameters as the configuration of the robot’s end effector.
Thus, the complete configuration of the object can be reformulated as $(x,y,\phi,\kappa(\cdot))$. Note that, up to this point, the description is exact and infinite dimensional - as $\kappa$ is a function. We move from infinite to finite domain by approximating the curvature with a Taylor expansion
$\kappa(s)\simeq\sum_{i=0}^{\infty}\theta_{i}s^{i},$ (3)
which can be truncated to degree $n$ to obtain a finite object configuration vector $\Theta\in\mathbb{R}^{n+1}$ approximating the real object to a desired precision. We use this to express the orientation of the central axis $\alpha$ (relative to the object base frame $\{S\}$) at a normalized distance $s\in[0,1]$ along the object’s length as the integral of the curvature function
$\alpha(s,\Theta)=\int_{0}^{s}\sum_{i=0}^{n}\theta_{i}v^{i}\mathrm{d}v$ (4)
The configuration vector for the floating base model combines the curvature and manipulator’s end effector variables: $q_{\mathrm{O}}=(\Theta,x,y,\phi)\in\mathbb{R}^{n+4}$.
Figure 2: Block diagram outline of the control strategy
II-C Object’s Full Kinematics Under Cosserat Assumption
So far, we have derived the object’s kinematic model under the implicit working assumption that it is infinitely thin. This can be readily relaxed under the Cosserat rod description [57] - i.e., we represent the object as nondeforming outside of its central backbone and assume that an undeformable slice of a point mass is rigidly connected at each point of the backbone.
Adding an offset of normalized distance $d\in[-\frac{1}{2},\frac{1}{2}]$ perpendicular to the central axis, we can then write the forward kinematics for a point on the object, scaling to its physical length $L$ and width $D$
$\begin{bmatrix}x_{\mathrm{B}}(s,d)\\ y_{\mathrm{B}}(s,d)\end{bmatrix}=\begin{bmatrix}x\\ y\end{bmatrix}+\begin{bmatrix}L\int_{0}^{s}\sin(\alpha_{\mathrm{B}}(s))\mathop{}\!\mathrm{d}v-Dd\cos(\alpha_{\mathrm{B}}(s))\\ L\int_{0}^{s}-\cos(\alpha_{\mathrm{B}}(s))\mathop{}\!\mathrm{d}v-Dd\sin(\alpha_{\mathrm{B}}(s))\end{bmatrix}.$ (5)
It is immediate to see that, for $d=0$, this expression reduces itself to the kinematics of the backbone.
II-D Object’s Dynamics
Now that the finite-dimensional forward kinematics describing the complete infinite-dimensional shape of the object is defined, we can proceed to establish the dynamics using the Euler-Lagrange methodology. These steps are only briefly summarized here, as they are very similar to the ones detailed in [40, 58]. The inertia matrix is constructed by consideration of the kinetic energy of each infinitesimal mass element in the object:
$B_{\mathrm{O}}(q_{\mathrm{O}})=\int_{0}^{1}\int_{-\frac{1}{2}}^{\frac{1}{2}}\rho(s,d)J_{s,d}^{T}J_{s,d}\,\mathop{}\!\mathrm{d}d\,\mathop{}\!\mathrm{d}s.$ (6)
Here $\rho(s,d)$ is the mass density distribution, and $J_{s,d}$ is the Jacobian matrix of the forward kinematics function (5) with respect to $q_{\mathrm{O}}$. The Coriolis and centrifugal matrix $C$ from the inertia matrix is evaluated by standard procedures - e.g., Christoffel symbols.
The gravitational force field is derived by differentiation of the gravitational potential energy of the infinitesimal masses, with the definition:
$G_{\mathrm{O}}(q_{\mathrm{O}})=g\nabla_{q_{\mathrm{O}}}\int_{0}^{1}\int_{-\frac{1}{2}}^{\frac{1}{2}}\rho(s,d)y_{\mathrm{B}}\,\mathop{}\!\mathrm{d}d\,\mathop{}\!\mathrm{d}s$ (7)
The direction of the gravitational field is considered directed along $-y$ in the $\{S_{\mathrm{B}}\}$ frame.
We model the object’s internal elastic and dissipative forces as discussed in [58]. An addition compared to that model is the relaxation of the assumption that the minimum of the elastic potential is in the straight configuration. Indeed, while this is a reasonable assumption for soft robots, it is not for deformable objects - as we will observe in the experimental validation of our results. This can be simply obtained by introducing an offset $\bar{\Theta}$ in the elastic force term - which is the configuration in which the object would rest when not immersed in a gravitational field.
Combining all these expressions yields the dynamic equations of motion for the floating base system:
$\small\begin{split}B_{\mathrm{O}}(q_{\mathrm{O}})\ddot{q}_{\mathrm{O}}&+C_{\mathrm{O}}(q_{\mathrm{O}},\dot{q}_{\mathrm{O}})\dot{q}_{\mathrm{O}}+G_{\mathrm{O}}(q_{\mathrm{O}})\\ +&\begin{bmatrix}kH&0_{n+1\times 3}\\ 0_{3\times n+1}&0_{3\times 3}\end{bmatrix}\begin{bmatrix}\Theta-\bar{\Theta}\\ x\\ y\\ \phi\end{bmatrix}\\ +&\begin{bmatrix}\beta H&0_{n+1\times 3}\\ 0_{3\times n+1}&0_{3\times 3}\end{bmatrix}\dot{q}_{\mathrm{O}}=\begin{bmatrix}0_{n\times 1}\\ F_{\mathrm{x}}\\ F_{\mathrm{y}}\\ \tau_{\phi}\end{bmatrix},\end{split}$ (8)
where $F_{\mathrm{x}}$, $F_{\mathrm{y}}$, and $\tau_{\phi}$ are generalized forces (two forces and torque) representing the action of the robot on the object, $H$ is the $n\times n$ Hankel matrix, and the internal stiffness and damping coefficients of the object averaged along its length are $k$ and $\beta$.
II-E Complete Dynamical Model
An object dynamics in the form expressed by (8) is extremely convenient when - as for our experimental setup - a force/torque sensor collocated at the gripper location is available that can directly measure $F_{\mathrm{x}},F_{\mathrm{y}},\tau_{\phi}$.
However, we want to be able to design controllers and assess stability when considering the complete robot-object system dynamics. First, we express the object’s forward kinematics as a function of the robot configuration $q_{\mathrm{r}}\in\mathbb{R}^{n_{\mathrm{r}}}$. This is achieved by plugging $(x,y,\phi)=h_{\mathrm{r}}(q_{\mathrm{r}})$ in (5) with $h_{\mathrm{r}}:\mathbb{R}^{n_{\mathrm{r}}}\rightarrow\mathrm{SO}(3)$ being the forward kinematics of the robot’s end effector.
Then, the full-order dynamics is readily derived by differentiating separately w.r.t. $q_{\mathrm{r}}$ and $\Theta$ the new kinetic and potential energies so defined
$\small\begin{split}&\begin{bmatrix}B_{\mathrm{r},\mathrm{r}}(q_{\mathrm{r}},\Theta)&B_{\mathrm{r},\Theta}(q_{\mathrm{r}},\Theta)\\ B_{\mathrm{r},\Theta}^{T}(q_{\mathrm{r}},\Theta)&B_{\Theta,\Theta}(\Theta)\end{bmatrix}\begin{bmatrix}\ddot{q}_{\mathrm{r}}\\ \ddot{\Theta}\end{bmatrix}+C(q_{\mathrm{r}},\Theta,\dot{q}_{\mathrm{r}},\dot{\Theta})\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}\\ &+\begin{bmatrix}0_{n_{\mathrm{r}}\times 3}&0_{n+1\times n_{\mathrm{r}}}\\ 0_{n_{\mathrm{r}}\times n+1}&\beta H\end{bmatrix}\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}+\begin{bmatrix}G_{\mathrm{r}}(q_{\mathrm{r}},\Theta)\\ G_{\Theta}(q_{\mathrm{r}},\Theta)\end{bmatrix}\\ &+\begin{bmatrix}0_{n_{\mathrm{r}}\times 3}&0_{n+1\times n_{\mathrm{r}}}\\ 0_{n_{\mathrm{r}}\times n+1}&kH\end{bmatrix}\begin{bmatrix}{q}_{\mathrm{r}}\\ (\Theta-\bar{\Theta})\end{bmatrix}=\begin{bmatrix}\tau\\ 0_{n+1\times 1}\end{bmatrix}.\end{split}$ (9)
where $(q_{\mathrm{r}},\Theta)\in\mathbb{R}^{n_{\mathrm{r}}+n+1}$ is the full order counterpart of $q_{\mathrm{O}}$ and complete configuration of the robot-object system. The other elements of the equation have their meaning in accordance with previous definitions.
Two facts about (9) are worth stressing here. First, the object and the robot are coupled purely through the potential field and the inertia forces. No elastic or dissipative coupling between the two dynamics is present.
Second, writing the operational space dynamics of the robot from (9) and equating it to (8) allows us to prove that
$(q_{\mathrm{r}},\dot{q}_{\mathrm{r}})\equiv 0\Rightarrow(F_{\mathrm{x}},F_{\mathrm{y}},\tau_{\phi})=\begin{bmatrix}0_{3\times n+1}&I_{3\times 3}\end{bmatrix}G_{\mathrm{O}}.$ (10)
III Manipulation as Closed-loop Control
The general control architecture proposed in this work is presented in Fig. 2. In the following, we present its main components step by step. In the process, we discuss the stability and convergence properties of a broad class of controllers that can be applied to this setting.
III-A Control Goal
Thanks to the effort to establish a complete model of the robot-object pair in Sec. II, we can now define our manipulation goal as a shape regulation task. More precisely, we wish to define a policy that specifies the control action $\tau$ in (9) such that
$\lim_{t\rightarrow\infty}\zeta(x,y,\phi,\Theta)=\zeta^{*}.$ (11)
Here, $\zeta:\mathbb{R}^{n+4}\rightarrow\mathbb{R}^{m}$ is a generic representation of a object-centric task. In other words, the configuration of the robot affects the fulfillment of the task only to the extent to which a change of $q_{\mathrm{r}}$ reflects a change in the robot’s state. This can, for example, be the position and orientation of one or more points along the object structure (we consider this case in the experimental validation), their relative position to an external target (e.g., insertion task), or directly the object shape $\Theta$.
III-B Low-level Control
Let’s consider the case in which the control inputs $\tau$ are used to regulate the collocated configuration variables $q_{\mathrm{r}}$.
III-B1 Stability of the zero dynamics
Combining (8) and (10) yields the zero dynamics
$\begin{split}B_{\Theta,\Theta}(\Theta)\ddot{\Theta}+C_{\Theta,\Theta}(\Theta,\dot{\Theta})\dot{\Theta}+\beta H\dot{\Theta}&+G_{\Theta}(\Theta,\phi^{*})\\ &+kH{(\Theta-\bar{\Theta})}=0.\end{split}$ (12)
Where $\phi^{*}$ is the end effector orientation associated with the fixed robot configuration $q^{*}_{\mathrm{r}}$. This collocated zero dynamics is such that
$\lim_{t\rightarrow\infty}\dot{\Theta}=0$ (13)
if $\beta>0$. Moreover, the ${\Theta}^{*}$ such that
$G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}=0,$ (14)
is an asymptotically stable equilibrium of the zero dynamics if the potential energy is convex in $\Theta^{*}$ (i.e., $k$ large enough). The stability is global if the potential energy is globally convex.
To prove these statements it is sufficient to take the object’s total energy as a Lyapunov candidate and follow similar steps as in the soft robotic survey paper [28], that we do not report here for the sake of space.
III-B2 Control strategy
The stability of the zero dynamics enables direct feedback linearization to regulate the robot’s states. However, we prefer here a simpler regulator that extends controllers commonly used in industrial robots.
$\tau=K_{\mathrm{P}}(q^{*}_{\mathrm{r}}-q_{\mathrm{r}})+K_{\mathrm{D}}(-\dot{q}_{\mathrm{r}})+G_{\mathrm{r}}(q^{*}_{\mathrm{r}},\Theta),$ (15)
where $K_{\mathrm{P}},K_{\mathrm{D}}$ are two positive definite control gain matrices. The only condition necessary for the closed loop stability is that $\beta>0$. No condition on $k$ is required. Again, for the sake of conciseness, we do not report the stability proof that can be obtained by recasting the steps of the one we propose in the soft robotic control paper [59].
Remarkably, this simple control strategy translates into the standard PD with gravity compensation when imposing the following hypothesis
$\frac{\partial G_{\mathrm{r}}}{\partial\Theta}\simeq 0,$ (16)
which is often verified in practice as it reflects the common case of the mass of the object being considerably smaller than the robot’s mass.
III-C High-level Control
Employing the low-level controller (15) allows to freely specify a configuration $q^{*}_{\mathrm{r}}$ for the robot. Combining this capability with standard Jacobian-based Cartesian controllers, we can use this capability to freely specify a configuration of the robot’s gripper $(x^{*},y^{*},\phi^{*})$.
To avoid destabilizing the low-level control loop, we cast the high-level component as pure feedforward, as outlined in the block diagram of Fig. 2. We propose a model-constrained nonlinear optimization as an object pose generation algorithm. We use the manipulator’s grasping point to control the equilibrium shape $\Theta^{*}$ of the object via the steady state zero dynamics equation (14). Note that we can do so because of the pure dependency of $G_{\Theta}$ on $(x^{*},y^{*},\phi^{*})$ rather than on the complete configuration $q^{*}_{\mathrm{r}}$.
Our desired object configuration is one such that the control goal (11) is verified. To find such a configuration, we define a cost based on a distance measure between desired and modeled positions and orientations, $d(\zeta^{*},\zeta)$.
Thus, the optimization problem can then be written as:
$\begin{split}&\min_{\Theta^{*}\in\mathbb{R}^{n+1},(x^{*},y^{*},\phi^{*})\in\mathbb{F}}\quad d(\zeta^{*},\zeta(x^{*},y^{*},\phi^{*},\Theta^{*}))\\ &\mathrm{s.t.}\quad G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}=0.\end{split}$ (17)
Here, we also impose a feasible set $\mathbb{F}$ on the floating base coordinates to represent the constrained reachable workspace of the manipulator - thus avoiding the risk of generating unfeasible references to the Cartesian control. This set can also be used to represent environmental constraints.
By solving (17), we retrieve a base configuration $(x^{*},y^{*},\phi^{*})$ minimizing the difference between the desired and modeled object shape, thus completing the control architecture in Fig. 2.
IV Experimental Validation of the Model
IV-A Experimental Setup
IV-A1 Test Objects
The DLOs used in our experimental investigations are heavy gauge electrical cables, chosen particularly due to their significant inertia and moderate stiffness. Different lengths, gauges (diameter and material composition) and the weight of the endpoint mass provide variation in the object properties leading to a range of different dynamic and equilibrium behaviours. Objects OB1-OB4 are made from the same, thicker cable material, and have variations of both a short or long length, and a significant or minimal endpoint mass. Objects OB5 and OB6 are made from a more flexible material, and only a long, weighted variation and a short, unweighted variation are considered. The full set of objects is pictured in Fig. 3.
TABLE I: Object Parameter Variations
ID $m_{L}[kg]$ $m_{1}[kg]$ L [m] D [m]
OB1 0.6 0.23 0.6 0.02
OB2 0.6 0.03 0.6 0.02
OB3 0.42 0.23 0.4 0.02
OB4 0.42 0.03 0.4 0.02
OB5 0.4 0.23 0.75 0.015
OB6 0.25 0.03 0.45 0.015
Figure 3: The test objects OB1-OB6 from left to right.
The objects considered here are assumed to have constant density and cross-section along their length; however, instead of using the exact continuous distribution of $\rho(s,d)$ when evaluating (6) and (7), we simplify using an approximation of six discrete lump masses. These are equally spaced along the length of the object at positions $s=\{\frac{1}{12},\frac{3}{12},\frac{5}{12},\frac{7}{12},\frac{9}{12},\frac{11}{12}\}$, each with a mass of $\frac{m_{L}}{6}$ distributed across the total diameter. Different masses $m_{0}$ and $m_{1}$ are also included at $s=0$ and $s=1$, to account for separate masses concentrated at the base and end of the object. In terms of unmodeled non-idealities, approximately 0.05-0.1m at the end of the objects is rigid to facilitate the attachment of the endpoint mass. Table I summarises the properties of the objects.
IV-A2 Test Equipment
A Franka Emika FR3 7-DoF manipulator arm was used to provide actuation of the object’s base in the experiments. When required, a Bota Systems SenseOne 6-axis force-torque sensor was assembled in between to measure the actuation applied. An Intel Realsense D345 camera was used to capture RGB images at 30Hz, from which the object state could be extracted. All hardware control and communication were managed through a PC running Ubuntu 20.04 and the ROS Noetic middleware. Figure 4 shows an overview of the test equipment setup.
Figure 4: The experimental setup used for validation of the proposed model and control architecture.
IV-B Extraction of the Object State
Points on the object at $s=\{0,0.5,1\}$ and $d=0$ were identified with coloured markers, facilitating the detection of their pixel locations via HSV thresholding. Under the assumption that the object stays confined to the plane on which the manipulator endpoint moves, the 3D marker positions were determined by intersecting this plane with the line projected through the associated pixel coordinate.
By taking the recovered 3D marker positions and discarding the third coordinate (in the direction normal to the assumed object plane) we are left with the planar model 2D Cartesian coordinates of the start-, mid- and endpoints of the central axis in $\{S_{B}\}$: $p_{s,B}$, $p_{m,B}$ and $p_{e,B}$. Note that hereafter we will omit the subscript $B$, so that any Cartesian coordinates mentioned can be assumed measured in $\{S_{B}\}$ unless indicated otherwise.
The floating base system configuration best matching the observed points on the object can be extracted by the numerical inverse kinematics algorithm 1, where $q_{\mathrm{O},init}$ is an initial guess, $p$ are the measured coordinates of the three marked points, and $h(q_{\mathrm{O}})$ and $J(q_{\mathrm{O}})^{+}$ evaluate the forward kinematics and Jacobian pseudoinverse respectively at the same locations. $\epsilon$ and $\Delta$ are simply tuned to a desired accuracy and convergence rate.
Algorithm 1 Numerical Inverse Kinematics
  $q_{\mathrm{O}}\leftarrow q_{\mathrm{O},init}$
  while $i\leq i_{max}$ do
     $e=\|p-h(q_{\mathrm{O}})\|_{2}$
     if $e<\epsilon$ then
        $q_{\mathrm{O}}\leftarrow q_{\mathrm{O}}$
        return  $q_{\mathrm{O}}$
     else
        $q_{\mathrm{O}}\leftarrow q_{\mathrm{O}}-\Delta J(q_{\mathrm{O}})^{+}e$
     end if
  end while
IV-C Parameter Identification
The model contains several object-specific parameters, some of which are not easy to directly measure. Measurable parameters are the object length $L$, diameter $D$, total mass of the object body $m_{L}$, object base mass $m_{0}$ and endpoint mass $m_{1}$. Immeasurable parameters are the object stiffness $k$, object damping $\beta$, and curvature offset $\bar{\Theta}$. To determine the unknown parameters, an identification procedure was developed in two parts, separating out those that could be determined from the static case and those that require dynamic analysis.
For the static case, the equilibrium state of the object is extracted at a series of discrete base orientation angles $\phi$. This data is used to construct a set of linear equations, from which a pseudoinverse solution can be used to determine the values for $k$ and $\bar{\Theta}$, providing a least squares best fit. Equation (18) shows the development of the zero dynamics in (14) into the form used to obtain the pseudoinverse solution. This represents a system of $n+1$ equations corresponding to a single $\phi$, in practice this is extended with additional rows for each $\phi$ and object state $\Theta^{*}$ measured, before the pseudoinversion step.
$\begin{split}G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}&=0\\ \begin{bmatrix}H\Theta^{*}&-H\end{bmatrix}\begin{bmatrix}k\\ k\bar{\Theta}\end{bmatrix}&=-G_{\Theta}(\Theta^{*},\phi^{*})\\ \begin{bmatrix}k\\ k\bar{\Theta}\end{bmatrix}&=\begin{bmatrix}H\Theta^{*}&-H\end{bmatrix}^{+}(-G_{\Theta}(\Theta^{*},\phi^{*}))\end{split}$ (18)
With the values of $k$ and $\bar{\Theta}$ estimated as outlined above, identification of the final unknown parameter $\beta$ can subsequently be done using data taken from a dynamic evolution of the object. The approach is similar, using the zero dynamics (12) to construct an overdetermined system of equations and calculate a pseudoinverse solution as shown in (19).
$\begin{split}\beta=(H\dot{\Theta})^{+}(-B_{\Theta,\Theta}(\Theta)\ddot{\Theta}-C_{\Theta,\Theta}(\Theta,\dot{\Theta})\dot{\Theta}&-G_{\Theta}(\Theta,\phi^{*})\\ &-kH(\Theta-\bar{\Theta})).\end{split}$ (19)
IV-D Model Validation Results
A polynomial of degree $n=1$ was used for the experimental validation: an affine curvature strain paramerisation [60] of the object shape, where $\Theta=(\theta_{0},\theta_{1})$. The parameter identification procedures laid out in Section IV-C were carried out using this model. The ultimately determined values are summarised in Table II, while the rest of this subsection will detail the practical considerations and intermediate results.
TABLE II: Identified Object Parameters
ID $k$ $\bar{\theta_{0}}$ $\bar{\theta_{1}}$ $\beta$
OB1 0.197 0.142 1.24 0.0347
OB2 0.226 -0.0358 2.06 0.0311
OB3 0.253 0.830 -0.452 0.0359
OB4 0.360 0.748 0.0711 0.0454
OB5 0.0761 0.354 0.636 0.0343
OB6 0.269 0.479 0.621 0.0397
IV-D1 Steady State
Static equilibrium datasets were collected for each of the objects at 23 orientations of $\phi$, angled at evenly spaced $\frac{\pi}{12}$ rad increments excluding $-\pi$. To minimize the influence of hysteresis on the object’s shape between measurements, they were returned to $\phi=0$ and shaken loose before moving to each measurement orientation. Configurations of $\Theta$ corresponding to the measured $p_{s}$, $p_{m}$ and $p_{e}$ were then extracted to use as input to the parameter identification formulation for $k$ and $\bar{\Theta}$ as described in Section IV-C. Note that the affine curvature model exhibits parameter-dependent multi-stability [61]. Here, it is generally safe to assume that the object will settle to an expected equilibrium if it begins in a nominal static state at $\phi=0$ and $\Theta$ close to $(0,0)$, and is rotated slowly within $-\pi\leq\phi<\pi$.
Following identification for the static case, we check how well the resulting model predicts the equilibrium behaviour of the real objects. Fig. 5a shows the results of this process for OB1, plotted for $\phi=\{0,\pm\frac{\pi}{4},\pm\frac{\pi}{2},\pm\frac{3\pi}{4}\}$ (7 of the 23 datapoints used). Here, the coloured crosses indicate the measured positions of $p_{m}$ and $p_{e}$, and the light orange curves represent the configurations of $\Theta$ from inverse kinematics, used in (18) to identify the parameters. The dark orange curves are the steady state of the model obtained by forward simulation using the identified parameters. A composite image of the actual object shapes at these angles is shown in Fig. 5b. The distance between the measured and modeled $p_{e}$ is plotted against the angle $\phi$ in Fig. 6 for the same object, providing a measure of the modeling accuracy.
Figure 5: a) Comparison of OB1 shape with $\Theta$ extracted from measured equilibria (light orange), and from modeling after parameter identification with the same data (dark orange). b) Composite image showing several of the measured equilibria for OB1.
Figure 6: Error between measured and modeled equilibrium endpoint at each angle $\phi$ in the steady state model validation for OB1.
Equivalent figures for the other objects are not included for the sake of space, but are available here. The results show that the static equilibrium model provides quite an accurate estimate for the endpoint position in the range $\frac{-3\pi}{4}<\phi<\frac{3\pi}{4}$ for all object types, with an error below 0.05m in almost all cases. Up to $\phi=\pm\pi$, most types continue to only suffer relatively mild deterioration in accuracy, with some notable exceptions. Most egregiously, OB4 and OB6 display a sudden spike in error for $\phi<\frac{-3\pi}{4}$; on closer inspection, the model with these properties is no longer multi-stable at this point, with the sole equilibrium point biased to the right-hand side of the $xy$ plane due to the offset curvature. Another limitation of the model is highlighted by OB5, with its worst-case combination of length, stiffness, and endpoint mass resulting in high curvature near the base followed by straightening out along the length. The affine curvature model cannot describe this over a single segment.
IV-D2 Dynamic Response
The data for dynamic identification was collected from a single evolution of a pendulum motion for each object with the base fixed. Fig. 7a shows a time lapse of the images captured for OB1, along with the detected positions of $p_{s}$, $p_{m}$ and $p_{e}$, as well as the $\Theta$ configurations from inverse kinematics projected into the camera frame. Note that this shows only the first swing, the full sets of data run until the oscillations cease.
This data was used to estimate the value of $\beta$, and the resulting fully identified model was then simulated from the same initial conditions. A plot comparing the evolution of $\Theta$ measured from the experiment, and generated by the simulation, is shown in Fig. 7b.
In Fig. 7c this simulated evolution is then projected back onto the same captured images to show more clearly how the simulated object shape compares to the real object in Cartesian space. Only every second frame is included for clarity, and it can be seen that after release the simulated object shape leads the actual motion of the object, which matches the phase difference visible in Fig. 7b.
Figure 7: a) Time lapse of partial extracted dynamic evolution of $\Theta$ for OB1. b) Comparison of the measured and simulated evolutions of the $\Theta$ states. c) Simulated evolution of $\Theta$ projected over captured images.
Equivalent plots to Fig. 7b for each object are included here, generally showing similarly closely matching dynamic responses, though the simulation tends to lead the real data for most object types. The time lapse images were not created for each object, but videos are available here.
V Experimental Validation of the Controller
The control framework was experimentally validated, focussing on control of the object endpoint as a practical goal specification.
V-A Position
In the first case we look at positioning the object endpoint at a goal $p^{*}$ only, so that $\zeta=p_{e}$, and take the Euclidean distance error as the optimization cost. We use a circular constraint on the floating base coordinates to represent the workspace of the manipulator, specifically a circle of radius 0.5m centred at $(0,0.333)$ in $\{S_{B}\}$, which approximates the reach of the FR3. Since the modeling accuracy and planar assumption start to deteriorate at higher base orientations, it was also decided to limit the range of angles to $\frac{-3\pi}{4}\leq\phi\leq\frac{3\pi}{4}$.
The optimization problem is therefore:
$\begin{split}&\min_{\Theta^{*}\in\mathbb{R}^{n+1},(x^{*},y^{*},\phi^{*})\in\mathbb{F}}\quad\|p^{*}-p_{e}(x^{*},y^{*},\phi^{*},\Theta^{*})\|_{2}\\ &\mathrm{s.t.}\quad G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}=0\\ \mathbb{F}=\{x^{2}&+(y-0.333)^{2}-0.5^{2}\leq 0\}\times\{\frac{-3\pi}{4}\leq\phi\leq\frac{3\pi}{4}\}.\end{split}$ (20)
We used the controller to position the object endpoint by solving (20) at each of a grid of endpoint goals, with 0.1m spacing. The grid covers an area in $\{S_{B}\}$ of $-0.7\leq x\leq 0.7$ m and $0.05\leq y\leq 0.75$ m, although depending on the object type it may not be necessary to cover the full height in the $y$-direction once it exceeds the reachable range. The MATLAB fmincon solver was used to obtain the solutions, and results were obtained for OB1, OB2, OB3 and OB5, as their properties cover a wide range of high deformation behaviour. Additionally, a ’model-free’ reference was evaluated, where $\phi=0$ and the manipulator is positioned at a constant offset to $p^{*}$ (or as close as possible given the manipulator workspace constraint), obtained from the measured steady state $p_{e}$.
Fig. 8 shows the manipulator solutions, and corresponding equilibrium configurations of the object model, used for one row of the workspace validation in both cases for OB1. Lines are drawn pairing the goals with the achieved modeled endpoints, and the red circle indicates the constraint placed on the manipulator endpoint position.
Figure 8: Solutions for one row of the endpoint positioning experiment using OB1, for the model-free reference case (top) and model-based case (bottom).
Based on these solutions, we assess the controller experimentally by measuring $p_{e}$ after moving the manipulator to each configuration, as demonstrated in the example image sequences in Figure 9. The results for OB1 are summarised in Figure 10a, where the grid colours correspond to the error between the goal $p^{*}$, located at the centre of each cell, and $p_{e}$. In the upper row of plots, the error is calculated to the modeled $p_{e}$ determined in the optimization, plotted as white dots; lines are drawn between these and the goals. In the lower row of plots, the error is calculated to the measured $p_{e}$, plotted as white dots; here the lines are drawn between the modeled and measured $p_{e}$.
Figure 9: Example image sequences from the endpoint position control validation experiment, with overlay of modeled final object goal states. a) OB1, b) OB2, c) OB3, d) OB5.
Figure 10: a) Visualisation of endpoint error over the workspace for OB1. Error from goal to modeled endpoint (upper row) and error from goal to measured endpoint (lower row). Model-free reference case (left) and model-based case (right). b) Mean endpoint positioning error taken over the full workspace grid, summarised for all tested objects.
Complete results for the four objects, including the solutions for each row and heatmap grid summaries are included here. For all objects it is clear that the endpoint reachability is significantly expanded compared to the reference case, as expected. The measured positioning accuracy in this increased workspace varies significantly, generally corresponding to decreased accuracy with the use of larger $|\phi|$. For further quantitative analysis, Fig. 10b plots the mean endpoint error over all measured grid points for each tested object, although it is noted that these values vary depending on the chosen area of grid coverage. Improvement in the model-based case is clearly demonstrated here, there is a reduction in the measured mean error compared to the reference case of 63.0%, 68.1%, 78.5% and 57.7% for the four objects respectively.
V-B Orientation
Here we continue assessing the reachable workspace of the object’s endpoint, now focussing only on the orientation, ie. the accuracy of positioning $p_{e}$ at a single location in $\{S_{B}\}$ but with the tip pointed in a specified angle. We will refer to the angle of the tip as $\psi_{e,B^{\prime}}=\phi+\alpha_{e}$, note that this is in the $\{S_{B^{\prime}}\}$ frame so that pointing directly downwards corresponds to 0. For this validation we only investigate OB4 and OB6, as inspection of the static identification results indicated a very limited achievable tip angle range for the other objects, due to their length and weight.
The experimental process is similar to that in Section V-A, now with $\zeta=(p_{e},\psi_{e,B^{\prime}})$ requiring the addition of the endpoint orientation error into the optimization cost. We simplify by ensuring the endpoint position goal $p^{*}$ is always exactly reachable by removing workspace position constraints on the manipulator, and only target endpoint orientation goals $\psi_{B^{\prime}}^{*}$ away from ±$\pm\pi$, so it is sufficient to define $d$ as the 2-norm of the combined position and angle errors. We relax the constraint on the manipulator angle to $\frac{-3\pi}{4}\leq\phi<\pi$, which allows a goal range of $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$. Hence the optimization problem here is:
$\begin{split}\min_{\Theta^{*}\in\mathbb{R}^{n+1},(x^{*},y^{*},\phi^{*})\in\mathbb{F}}\quad&\begin{Vmatrix}p^{*}-p_{e}(x^{*},y^{*},\phi^{*},\Theta^{*})\\ \psi_{B^{\prime}}^{*}-\psi_{e,B^{\prime}}(\phi^{*},\Theta^{*})\end{Vmatrix}_{2}\\ \mathrm{s.t.}\quad G_{\Theta}(\Theta^{*},\phi^{*})+&kH{(\Theta^{*}-\bar{\Theta})}=0\\ \mathbb{F}=\{\frac{-3\pi}{4}&\leq\phi\leq\pi\}.\end{split}$ (21)
Solving the problem for $\psi_{B^{\prime}}^{*}\in[\frac{-\pi}{2},\frac{\pi}{2}]$ at $\frac{\pi}{12}$ rad increments and $p^{*}=(0,0.4)$ produced the configurations shown in Fig. 11a for OB4. Fig. 11b displays a composite of the real object state at each of these configurations.
To quantitatively assess the accuracy, the angle of the tip was determined by measuring the points at the start and end of the endpoint weight to obtain a direction vector in the $xy$ plane. The error between this and the goal angle is plotted in Fig. 12. Along with the results for OB6, available here, this investigation also reveals considerable ability to reach specified orientations in the range $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$, with a recorded maximum absolute angle error of around 0.29 rad or 17°. Particularly in the range of $0\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$ the error is maintained below 0.12 rad (7°) and 0.21 rad (12°) for OB4 and OB6 respectively.
Figure 11: a) Solutions for the endpoint orientation control experiment with OB4. b) Composite image showing real object configurations in the experiment. The cross indicates the intended endpoint position.
Figure 12: Measured tip angle error in the orientation control experiment with OB4.
VI Conclusion
In this work, we have presented a novel approach to the manipulation of deformable linear objects using a strain-based modeling approach. The choice of a polynomial curvature parametrization allows for a compact manipulator-object system state that is suitable for model-based control. We then formally define the manipulation goal as a control problem, for which we introduce control strategies with provable guarantees. Our experimental results show that in this form, a model of minimal dimension is accurate enough for regulation of the steady-state endpoint position and orientation over a range of object properties. Our results demonstrate average positioning accuracy improvements of between 57.7% and 78.5% over a model-free reference approach and maximum orientation errors of 12°and 17°. We also show results assessing the accuracy of the full dynamic model. Future work will focus on closing the loop on the object state (non-collocated control) and performing more dynamic motion, starting from trajectory tracking and leading to accurate throwing and whipping.